OpenStudy (anonymous):

john has a balance of $3,000 on his credit card that charges 1% interest per month on any unpaid balance. John can afford to pay $100 toward the balance each month. His balance after making a $100 payment is given by the recursively defined sequence B of o= $3,000 B of n=1.01B of (n-1)-100 (a) determine john's balance after making the first payment. that is, determine B of 1 (b) using a graphing utility, determine when john's balance will be below $2,000. How many payments of $100 have been made? (c) using a graphing utility, determine when john will pay off the balance. what is the total of

OpenStudy (anonymous):

all the payments? (d) what was john's interest expense?

OpenStudy (anonymous):

:0

OpenStudy (anonymous):

hahaah yeah its a tough one. don't know where to start

OpenStudy (a_clan):

(a) First part is to solve a recursive equation Bof( n ) =1.01* Bof(n-1) - 100 ....(If I am getting the formula right) So, Bof( 1 ) = 1.01*Bof( 0 ) -100 We know, Bof(0)= 3000 Substitute the value, Bof(1) =1.01*3000 -100 ......Solving this will give the answer